A semicircular plate of mass 'm' has radius ‘r’ and centre 'c'. The centre of mass of the plate is at a distance 'x' from its centre 'c'. Its moment of inertia about an axis passing through its centre of mass and perpendicular to its plane is
1
Expert's answer
2019-08-23T09:34:35-0400

Moment of inertia of a semi circular plate about its centre C="\frac{mr^2}{2}"

Moment of inertia about axis passing through a distance "x" from centre=

"I_x=I_c-mx^2"

"I_x=\frac{mr^2}{2}-mx^2"

This x is the centre of mass distance which is equal to "\frac{4r}{3 \pi}"

Substituting this value of x,

we get,

"I_x=\frac{mr^2}{2}-m{(\frac{4r}{3 \pi})}^2"

"I_x=\frac{mr^2}{2}-\frac{16 m r^2}{9 \pi ^2}"